OpenStudy (anonymous):

Hannah wishes to find the vertex of the function x2 + 3x - 28 = 0 by completing the square. Which of the following equations may be used to identify the coordinates of the vertex?

OpenStudy (anonymous):

A. (x + 1.2)2 = 30.25 B. (x + 2.25)2 = 30.5 C. (x + 2.5)2 = 30.75 D. (x + 2.75)2 = 32.5

OpenStudy (anonymous):

@sweetburger

OpenStudy (anonymous):

@blanderos24

OpenStudy (anonymous):

iven y=x^2+3x-28. How do I algebraically find: The vertex, The x and y intercepts, and write the equation in vertex and factor form. ---- Complete the square: y = x^2+3x + 9 -9-28 --- y = (x+3)^2-37 ---- Vertex: (-3,-37) ------------------------ y-intercept: Use the form y = x^2+3x-28